OpenStudy (anonymous):

Solve the equations: (x-1)^2+(y+2)^2=8, x^2+y^2-3x+y=12. ANSWER {(3,-4),(-9/5,-12/5)}

OpenStudy (anonymous):

subtract first from the second and find y in terms of x

OpenStudy (anonymous):

i think i have 2 simplify the 2 equations first then subtract them

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

(x-1)^2+(y+2)^2=8 x^2-2x+1+y^2+4y+4=8 x^2+y^2-2x+4y-3=0---------( equation 1)

OpenStudy (anonymous):

x^2+y^2-3x+y=12 x^2+y^2-3x+y-12=0---------(equation 2)

OpenStudy (anonymous):

am i doing it right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Subtraction